### v1

# Toric varieties modulo reflections

Abstract.
Let $W$ be a finite group generated by reflections of a lattice $M$. If a lattice polytope $P\subset M\otimes_{\mathbb{Z}}\mathbb{R}$ is preserved by $W$, then we show that the quotient of the projective toric variety $X_{P}$ by $W$ is isomorphic to the toric variety $X_{P\cap D}$, where $D$ is a fundamental domain for the action of $W$. This answers a question of Horiguchi-Masuda-Shareshian-Song, and recovers results of Blume, of the second author, and of Gui-Hu-Liu.
1. Introduction
1.1.
The quotient of a toric variety by a finite group need not be toric, even when the group acts by toric automorphisms. For instance, the quotient of $\mathbb{P}^{n}\times\mathbb{P}^{n}$ by the $\mathbb{Z}/2$ action that swaps the factors is not toric when $n\geq 2$ [HMJJ21, Remark 8.2], and the quotient of a 47-torus over $\mathbb{Q}$ by the $\mathbb{Z}/47$ action that cycles its factors is not even rational [S69, Theorem 1] [^1] We thank Laurent Moret-Bailly for writing the MathOverflow post that made us aware of this example.! The existence of these examples is unsuprising: it is almost never possible to make the quotient map torus-equivariant, so there is little reason to expect a toric structure on the quotient.
[…]
$s_{\alpha}(v):=v-2\frac{\left\langle v,\alpha\right\rangle}{\left\langle\alpha,\alpha\right\rangle}\alpha$
for some nonzero $\alpha\in V$. A reflection group in $V$ is a finite subgroup of $\operatorname{GL}(V)$ generated by reflections. The fundamental domain for the action of a finite reflection group $W$ in $V$ is a polyhedral cone $D\subset V$.
Theorem 1.1.
[L05, Theorem 6.1.1] If $W\subset\operatorname{GL}(V)$ is a finite reflection group preserving a lattice $M\subset V$, then there is an isomorphism $\mathbb{Z}[D\cap M]\to\mathbb{Z}[M]^{W}$.
The semigroup $D\cap M$ is an example of a saturated affine semigroup, i.e. a finitely generated $\mathbb{Z}_{\geq 0}$-submodule $S\subset M$ such that for all $k\in\mathbb{Z}_{\geq 0}$ and $m\in M$, $km\in S$ implies $m\in S$. Saturated affine semigroup algebras are precisely the the coordinate rings of normal affine toric varieties [CLS11, Theorem 1.3.5], so we may restate Theorem 1.1 as:
Theorem 1.1.
[L05, Theorem 6.1.1] If $W\subset\operatorname{GL}(V)$ is a reflection group preserving a lattice $M\subset V$, and $T=\operatorname{Spec}\mathbb{Z}[M]$, then there is an isomorphism of algebraic varieties $T/W\cong\operatorname{Spec}\mathbb{Z}[D\cap M]$.
We emphasize that the quotient map is essentially never toric.
1.2.
The main result of this note is a generalization of Lorenz’s theorem.
Theorem 1.2.
If $W\subset\operatorname{GL}(V)$ is a reflection group preserving a lattice $M\subset V$ and a saturated affine semigroup $S\subset M$, then there is an isomorphism $\mathbb{Z}[D\cap S]\to\mathbb{Z}[S]^{W}$.
Equivalently, we may state Theorem 1.2 as:
Theorem 1.2.
If $W\subset\operatorname{GL}(V)$ is a reflection group preserving a lattice $M\subset V$ and a saturated affine semigroup $S\subset M$, and $X=\operatorname{Spec}\mathbb{Z}[S]$, then $X/W$ is isomorphic to the normal affine toric variety $\operatorname{Spec}\mathbb{Z}[D\cap S]$.
Projective toric varieties with dense torus $\operatorname{Spec}\mathbb{Z}[M]$ correspond to full-dimensional lattice polytopes $P\subset V$ [CLS11, Chapter 2]. A consequence of Theorem 1.2 is the following statement, conjectured in [HMJJ21, Question 8.1].
Corollary 1.3.
Let $W\subset\operatorname{GL}(V)$ be a reflection group preserving a lattice $M\subset V$ and a full-dimensional lattice polytope $P\subset V$, with associated projective toric variety $X_{P}$. The quotient $X_{P}/W$ is isomorphic to the projective toric variety $X_{P\cap D}$.
Special cases of Corollary 1.3 include the well-known isomorphisms $(\mathbb{P}^{1})^{n}/S_{n}\cong\mathbb{P}^{n}$ and $\mathbb{P}^{n-1}/S_{n}\cong\mathbb{P}(1,2,\ldots,n)$, as well as results of Blume [B15, Theorems 3.1, 5.3, 5.9], which state that the projective toric variety associated to a permutohedron of type A, B, or C modulo the corresponding Weyl group is toric. Corollary 1.3 also strengthens the main theorem of the second author [G24], which says that the quotient of the toric variety of a dominant weight polytope by a parbolic subgroup of a Weyl group of appropriate type is homotopy equivalent to a toric variety. Passing to cohomology rings, Corollary 1.3 recovers the main theorem of [HMJJ21], as well as [GHL24, Theorems 1.2] for crystallographic reflection groups. Unlike [HMJJ21, GHL24], we do not produce an explicit isomorphism $H^{*}(X_{P\cap D})\to H^{*}(X_{P})^{W}$.
Future work
In forthcoming work, we will treat the case of arbitrary fans. In particular, we will describe the normal fan of the polytope $P\cap D$ and consider quotients of toric varieties that are neither affine nor projective.
Acknowledgements
We are grateful to Tao Gui for making us aware of [HMJJ21, Question 8.1]. We also thank Chiara Damiolini & Sam Payne for discussing early versions of this project. CC and CS thank the Simons Foundation for support; CC is also supported by NSF grant DMS-2039316.
2. Finite reflection groups
We review the theory of finite reflection groups, following [H90].
2.1.
[…]
$s_{\alpha}(v):=v-2\frac{\left\langle v,\alpha\right\rangle}{\left\langle\alpha,\alpha\right\rangle}\alpha.$
A reflection group in $V$ is a subgroup of $\operatorname{GL}(V)$ generated by reflections. All reflection groups appearing in this paper are assumed finite.
2.2.
A root system is a finite set $\Phi\subset V$ such that for all $\alpha\in\Phi$,
  1. (i)
    $\phi\cap\mathbb{R}\left\langle\alpha\right\rangle=\{\alpha,-\alpha\}$, and
  2. (ii)
    $s_{\alpha}(\Phi)=\Phi$.
Elements of $\Phi$ are roots. If $\Phi$ is a root system, then the subgroup $W$ of $\operatorname{GL}(V)$ generated by $\{s_{\alpha}:\alpha\in\Phi\}$ is a finite reflection group, and all such groups arise in this manner. Any root system contains a simple system, that is, is a linearly independent subset $\Delta\subset\Phi$ such that
  1. (i)
    $\Delta$ and $\Phi$ span the same subspace of $V$, and
  2. (ii)
    every element of $\Phi$ is real linear combination of elements of $\Delta$ with coefficients all of the same sign.
Elements of $\Delta$ are simple roots. Properties of $\Delta$ include:
  1. (i)
    $W$ is generated by $\{s_{\alpha}:\alpha\in\Delta\}$ [H90, Theorem 1.5].
  2. (ii)
[…]
Remark 2.1.
[…]
$2\left\langle\alpha,v\right\rangle=\left\langle\alpha,v\right\rangle+\left\langle s\alpha,sv\right\rangle=\left\langle\alpha,v\right\rangle+\left\langle-\alpha,v\right\rangle=0,$
so $s=s_{\alpha}$, which is a reflection as defined here.
Remark 2.2.
The definition of “root system” used here is that of [H90], but differs slightly from the definition used in some other sources. To assure the reader that this discrepancy causes no problem, we will provide proofs for results from references other than [H90].
2.3.
Fix a simple system $\Delta$ for $W$, with corresponding fundamental domain $D$. Define a partial order on $V$ by $u\geq v$ if and only if $u-v$ is a non-negative real linear combination of simple roots.
Proposition 2.3.
[H90, Lemma 1.12] If $v\in D$, then $v-wv\in\mathbb{Z}_{\geq 0}\left\langle\alpha_{1},\ldots,\alpha_{r}\right\rangle$ for all $w\in W$. In particular, $v\geq wv$.
Let $\operatorname{conv}(A)$ denote the convex hull of a subset $A\subset V$.
Proposition 2.4.
[BtD85, Proposition VI.2.4(ii)] If $u,v\in D$, then $u\geq v$ if and only if $\operatorname{conv}(Wu)$ contains $v$.
Proof.
If $v\in\operatorname{conv}(Wu)$, then there are non-negative real numbers $c_{w}$, $w\in W$, such that $\sum_{w}c_{w}=1$ and $v=\sum_{w}c_{w}wu$. By Proposition 2.3, $wu\leq u$, so
$v=\sum_{w}c_{w}wu\leq\sum_{w}c_{w}u=u.$
Conversely, suppose $v\not\in\operatorname{conv}(Wu)$. Then there is $l\in V$ such that for all $w\in W$, $\left\langle v,l\right\rangle>\left\langle wu,l\right\rangle=\left\langle u,w^{-1}l\right\rangle$. In particular, if $l^{\prime}=w^{-1}l\in D$, then by Proposition 2.3,
$\left\langle v,l^{\prime}\right\rangle\geq\left\langle w^{-1}v,l^{\prime}\right\rangle=\left\langle v,wl^{\prime}\right\rangle>\left\langle u,l^{\prime}\right\rangle,$
hence $\left\langle u-v,l^{\prime}\right\rangle<0$. This proves that $u\not\geq v$ because an element of $\mathbb{R}_{\geq 0}\left\langle\alpha_{1},\ldots,\alpha_{r}\right\rangle$ has non-negative inner product with all elements of $D$. ∎
2.4.
Call a finite reflection group $W$ crystallographic if it preserves a lattice in $V$ (i.e. the $\mathbb{Z}$-span of a basis of $V$). Given a simple system $\Delta=\{\alpha_{1},\ldots,\alpha_{r}\}$ for $W$, the associated weight group is
[…]
The weight group decomposes as
$\Lambda=Z\oplus\mathbb{Z}\left\langle\lambda_{1},\ldots,\lambda_{r}\right\rangle,$
where $Z=\cap_{i=1}^{r}\alpha_{i}^{\perp}$ and the vectors $\lambda_{1},\ldots,\lambda_{r}\in Z^{\perp}$ are the fundamental weights, determined by
$2\frac{\left\langle\lambda_{i},\alpha_{j}\right\rangle}{\left\langle\alpha_{j},\alpha_{j}\right\rangle}=\begin{cases}1,&i=j\\ 0,&\text{else.}\end{cases}$
The fundamental weights are linearly independent. Likewise, the fundamental domain decomposes as
$D=Z\times\mathbb{R}_{\geq 0}\left\langle\lambda_{1},\ldots,\lambda_{r}\right\rangle.$
Proposition 2.5.
[speyernotes, §15] Let $W$ be a finite reflection group that preserves a lattice $M\subset V$. There is a simple system $\Delta$ for $W$ such that
$\Delta\subset M\subset\Lambda.$
Morevoer, the action of $W$ preserves $\Lambda$.
Proof.
Let $\Delta_{0}$ be a simple system for $W$. If $\alpha\in\Delta_{0}$ and $m\in M\setminus\alpha^{\perp}$, then
$M\ni m-s_{\alpha}(m)=2\frac{\left\langle m,\alpha\right\rangle}{\left\langle\alpha,\alpha\right\rangle}\alpha,$
so $M\cap\mathbb{R}\left\langle\alpha\right\rangle\neq\{0\}$. Consequently, $M\cap\mathbb{R}\left\langle\alpha\right\rangle$ is a rank 1 lattice with a unique generator $\alpha^{\prime}$ satisfying $\left\langle\alpha,\alpha^{\prime}\right\rangle>0$. The set $\Delta=\{\alpha^{\prime}:\alpha\in\Delta_{0}\}$ is a simple system for $W$ that is contained in $M$.
The weight lattice $\Lambda$ associated to $\Delta$ contains $M$ because for any $\alpha\in\Delta$ and $m\in M$,
$\mathbb{Z}\left\langle\alpha\right\rangle=\mathbb{R}\left\langle\alpha\right\rangle\cap M\ni m-s_{\alpha}(m)=2\frac{\left\langle m,\alpha\right\rangle}{\left\langle\alpha,\alpha\right\rangle}\alpha.$
The action of $W$ preserves $\Lambda$ because for each $\alpha\in\Delta$ and $\lambda\in\Lambda$,
$s_{\alpha}(\lambda)=\lambda-2\frac{\left\langle\lambda,\alpha\right\rangle}{\left\langle\alpha,\alpha\right\rangle}\alpha\in\Lambda+\mathbb{Z}\left\langle\alpha\right\rangle=\Lambda,$
and the reflections associated to simple roots generate $W$. ∎
3. Proofs
Throughout this section, we fix notations:
[…]
  • $M\subset V$ is a lattice,
  • $W\subset\operatorname{GL}(V)$ is a finite reflection group preserving $M$, with simple system $\Delta=\{\alpha_{1},\ldots,\alpha_{r}\}\subset M$ as in Proposition 2.5,
  • $\lambda_{1},\ldots,\lambda_{r}$ the corresponding fundamental weights,
  • $D=Z\times\mathbb{R}_{\geq 0}\left\langle\lambda_{1},\ldots,\lambda_{r}\right\rangle$ the corresponding fundamental domain, and
  • $\Lambda=Z\oplus\mathbb{Z}\left\langle\lambda_{1},\ldots,\lambda_{r}\right\rangle\supset M$ the weight group, and $\Lambda_{\circ}=\mathbb{Z}\left\langle\lambda_{1},\ldots,\lambda_{r}\right\rangle$.
All future mention of “simple roots” will refer to elements of $\Delta$. If $S$ is a semigroup, let $\mathbb{Z}[S]$ denote its semigroup algebra, i.e. the group of finite $\mathbb{Z}$-linear combinations of symbols $\{\chi^{s}:s\in S\}$ with multiplication given by $\chi^{s}\chi^{t}=\chi^{s+t}$.
3.1.
The decompositions in Section 2.4 imply that
$\mathbb{Z}[D\cap\Lambda]$ $\cong\mathbb{Z}[Z]\otimes_{\mathbb{Z}}\mathbb{Z}[\Lambda_{\circ}]=\mathbb{Z}[Z]\otimes_{\mathbb{Z}}\mathbb{Z}[\chi^{\lambda_{1}},\ldots,\chi^{\lambda_{r}}],$
$\mathbb{Z}[\Lambda]^{W}$ $\cong\mathbb{Z}[Z]\otimes_{\mathbb{Z}}\mathbb{Z}[\Lambda_{\circ}]^{W}=\mathbb{Z}[\Lambda]^{W}\cong\mathbb{Z}[Z]\otimes_{\mathbb{Z}}\mathbb{Z}[\chi^{\lambda_{1}},\ldots,\chi^{\lambda_{r}}]^{W},$
and that the generators $\chi^{\lambda_{1}},\ldots,\chi^{\lambda_{r}}$ are algebraically independent. Consequently, there is a well-defined map of $\mathbb{Z}$-algebras
$\Psi_{\circ}:\mathbb{Z}[D\cap\Lambda_{\circ}]\to\mathbb{Z}[\Lambda_{\circ}]^{W}$
determined by
$\Psi_{\circ}(\chi^{\lambda_{i}})=\sum_{v\in W\lambda_{i}}\chi^{v},\quad 1\leq i\leq r.$
This map extends uniquely to a map of $\mathbb{Z}[Z]$-algebras
$\Psi:\mathbb{Z}[D\cap\Lambda]\to\mathbb{Z}[\Lambda]^{W}.$
Theorem 3.1.
[B02, Theorem VI.3.4.1] $\Psi_{\circ}$ is an isomorphism.
Theorem 3.1 implies $\Psi$ is also an isomorphism. Since our proof of Theorem 1.2 already requires analysis of $\Psi$, we provide a proof of Theorem 3.1 in Section 3.3.
3.2.
The elements
$\underline{\chi}^{u}=\sum_{v\in Wu}\chi^{v},\quad u\in D\cap\Lambda$
are a $\mathbb{Z}$-basis of $\mathbb{Z}[\Lambda]^{W}$, and the support of $f=\sum_{u\in D\cap\Lambda}a_{u}\underline{\chi}^{u}$ is
$\operatorname{supp}(f):=\{u\in D\cap\Lambda:a_{u}\neq 0\}.$
We study the supports of elements $\Psi(\chi^{u})$.
Lemma 3.2.
Let $u\in D\cap\Lambda_{\circ}$.
  1. (i)
    If $v\in\operatorname{supp}\Psi_{\circ}(\chi^{u})$, then $u-v\in\mathbb{Z}_{\geq 0}\left\langle\alpha_{1},\ldots,\alpha_{r}\right\rangle$. In particular, $u\geq v$.
  2. (ii)
    $\underline{\chi}^{u}$ has coefficient 1 in $\Psi_{\circ}(\chi^{u})$.
Proof.
Write $u=\sum_{i=1}^{r}b_{i}\lambda_{i}$ with $b_{1},\ldots,b_{r}\in\mathbb{Z}_{\geq 0}$. We may expand
$\Psi(\chi^{u})=\Psi_{\circ}(\chi^{u})=\prod_{i=1}^{r}\left(\sum_{v\in W\lambda_{i}}\chi^{v}\right)^{b_{i}}=\sum_{v\in D\cap\Lambda_{\circ}}a_{v}\underline{\chi}^{v},$
where
$a_{v}=\#\left\{(v_{ij})\in\prod_{\begin{subarray}{c}i=1,\ldots,r\\ j=1,\ldots,b_{i}\end{subarray}}W\lambda_{i}:\sum_{ij}v_{ij}=v\right\}.$
If $v\in\operatorname{supp}\Psi(\chi^{u})$, then
$v=\sum_{\begin{subarray}{c}i=1,\ldots,r\\ j=1,\ldots,b_{i}\end{subarray}}v_{ij}$
for some $(v_{ij})\in\prod_{\begin{subarray}{c}i=1,\ldots,r\\ j=1,\ldots,b_{i}\end{subarray}}W\lambda_{i}$. By Proposition 2.3, $v_{ij}\leq\lambda_{i}$ for all $i$, so
$v=\sum_{\begin{subarray}{c}i=1,\ldots,r\\ j=1,\ldots,b_{i}\end{subarray}}v_{ij}\leq\sum_{\begin{subarray}{c}i=1,\ldots,r\\ j=1,\ldots,b_{i}\end{subarray}}\lambda_{i}=u.$
This inequality is an equality if and only if $v_{ij}=\lambda_{i}$ for all $i,j$, so $a_{u}=1$. Finally, $u-v\in\mathbb{Z}\left\langle\alpha_{1},\ldots,\alpha_{r}\right\rangle$ because $\lambda_{i}-v_{ij}\in\mathbb{Z}\left\langle\alpha_{1},\ldots,\alpha_{r}\right\rangle$ for all $i,j$. ∎
Lemma 3.3.
If $S\subset M$ is a saturated affine semigroup preserved by the action of $W$, and $u\in D\cap S$, then $\operatorname{supp}\Psi(\chi^{u})\subset S$.
Proof.
Let $C:=\mathbb{R}_{\geq 0}\left\langle s:s\in S\right\rangle\subset V$. Plainly, $C$ is convex and preserved by $W$. Moreover, $S=C\cap M$ by [CLS11, Theorem 1.3.5].
Let $u\in D\cap S$ and $v\in\operatorname{supp}\Psi(\chi^{v})$. We may write uniquely $u=z+u_{\circ}$ and $v=z+v_{\circ}$, with $v_{\circ}\in\operatorname{supp}\Psi_{\circ}(\chi^{u_{\circ}})$.
We first show that $v\in M$. By Lemma 3.2, $u-v=u_{\circ}-v_{\circ}\in\mathbb{Z}_{\geq 0}\left\langle\alpha_{1},\ldots,\alpha_{r}\right\rangle$. Since $u,\alpha_{1},\ldots,\alpha_{r}\in M$, $v$ must also lie in $M$.
Next, we show that $v\in C$. By Lemma 3.2, $u_{\circ}\geq v_{\circ}$. By Proposition 2.4, this means $v_{\circ}$ is in the convex hull of $Wu_{\circ}$. Consequently,
$C\supset\operatorname{conv}(Wu)=z+\operatorname{conv}(Wu_{\circ})\ni z+v_{\circ}=v.$
We conclude that $v\in C\cap M=S$. ∎
3.3.
We now move to establish the main results.
Lemma 3.4.
Let $R$ be any commutative ring. If $U\subset D\cap\Lambda_{\circ}$ is a finite set of weights, then $\{\Psi_{\circ}(\chi^{u}):u\in U\}\subset R\otimes_{\mathbb{Z}}\mathbb{Z}[\Lambda_{\circ}]$ is linearly independent over $R$.
Proof.
We induct on the cardinality of $U$. If #$\#U=1$, the statement is true by Lemma 3.2ii.
Otherwise, if #$\#U>1$, let $v\in U$ be a maximal element with respect to $\leq$. By Lemma 3.2, there is no $v^{\prime}\in U\setminus v$ such that $v\in\operatorname{supp}\Psi_{\circ}(\chi^{v^{\prime}})$; however, $v\in\operatorname{supp}\Psi_{\circ}(\chi^{v})$. Consequently, $\Psi_{\circ}(\chi^{v})$ is not in the span of $\{\Psi_{\circ}(\chi^{u}):u\in U\setminus v\}$. By the induction hypothesis, it follows that $\{\Psi_{\circ}(\chi^{u}):u\in U\}$ is linearly independent. ∎
Proof of Theorem 3.1.
We first show $\Psi_{\circ}$ is injective. If $f\in\ker(\Psi_{\circ})$, then $f$ is contained in a subgroup of the form $\mathbb{Z}\left\langle U\right\rangle:=\mathbb{Z}\left\langle\chi^{u}:u\in U\right\rangle$ with $U\subset D\cap\Lambda_{\circ}$ a finite set. Lemma 3.4 implies that $\Psi_{\circ}$ is injective on $U$, so $f=0$.
Next, we show that $\Psi_{\circ}$ is surjective. To do this, we will induct to show that $\underline{\chi}^{u}$ is in the image of $\Psi_{\circ}$ for each $u\in D\cap\Lambda_{\circ}$. Given $u\in D\cap\Lambda_{\circ}$, let
$n(u):=\#\{v\in D\cap\Lambda_{\circ}:v\leq u\}.$
We have $n(u)=1$ if and only if $u$ is a minimal element of $D\cap\Lambda_{\circ}$ with respect to $\leq$ (minimal elements exist because for any $u\in D\cap\Lambda_{\circ}$, $\{v\in D\cap\Lambda_{\circ}:v\leq u\}$ is the intersection of the discrete set $\Lambda_{\circ}$ with a compact set by Proposition 2.4, hence finite).
If $n(u)=1$, then Lemma 3.2 implies $\Psi_{\circ}(\chi^{u})=\sum_{v\in Wu}\chi^{v}=\underline{\chi}^{u}$. Otherwise, suppose $n(u)>1$. By the induction hypothesis and Lemma 3.2,
$\Psi_{\circ}(\chi^{u})-\underline{\chi}^{u}\in\mathbb{Z}\left\langle\underline{\chi}^{v}:v<u\right\rangle\subset\mathbb{Z}\left\langle\underline{\chi}^{v}:n(v)<n(u)\right\rangle\subset\Psi_{\circ}(\mathbb{Z}[D\cap\Lambda_{\circ}]),$
so $\underline{\chi}^{u}$ is in the image of $\Psi_{\circ}$. This completes the proof. ∎
Proof of Theorem 1.2.
By Theorem 3.1, $\Psi_{\circ}$ is an isomorphism. Since $\mathbb{Z}[Z]$ is free over $\mathbb{Z}$, it follows that $\Psi$ is also an isomorphism. It now suffices to show that $\Psi(\mathbb{Z}[D\cap S])=\mathbb{Z}[S]^{W}$.
The containment $\Psi(\mathbb{Z}[D\cap S])\subset\mathbb{Z}[S]^{W}$ holds by Lemma 3.3. For the reverse containment, suppose $f\in\mathbb{Z}[D\cap\Lambda]$ is not in $\mathbb{Z}[D\cap S]$. There there is a maximal $v\in D\cap\Lambda$ such that $v\not\in S$, but $\chi^{v}$ appears in $f$. By Lemma 3.2 and Lemma 3.3, if $\chi^{v^{\prime}}$ appears in $f$ with $v^{\prime}\neq v$, then $v\not\in\operatorname{supp}\Psi(\chi^{v^{\prime}})$; therefore, $v\in\operatorname{supp}\Psi(f)$. It follows that $\Psi(f)$ is not contained in $\mathbb{Z}\left\langle\chi^{s}:s\in S\right\rangle$, hence is also not contained in
$\mathbb{Z}[S]^{W}=\mathbb{Z}\left\langle\chi^{s}:s\in S\right\rangle\cap\mathbb{Z}[\Lambda]^{W}.$
We conclude that $\Psi(f)\in\mathbb{Z}[S]^{W}$ if and only if $f\in\mathbb{Z}[D\cap S]$. ∎
Proof of Theorem 1.2.
In light of [CLS11, Theorem 1.3.5], we need only check that the linear functionals $f_{\alpha}=\left\langle\alpha,-\right\rangle$, $\alpha\in\Delta$, defining $D$ can be rescaled to lie in $\operatorname{Hom}(M,\mathbb{Z})$. This is possible because $M\subset\Lambda$, so $\frac{2f_{\alpha}}{\left\langle\alpha,\alpha\right\rangle}$ takes integer values on $M$. ∎
Proof of Corollary 1.3.
Recall from [CLS11, Chapter 2] that a lattice polytope $P\subset V$, gives rise to the projective toric variety $X_{P}:=\operatorname{Proj}\mathbb{Z}[tP]$, where $\mathbb{Z}[tP]$ is the saturated affine semigroup
$S:=\{(t,p):t\in\mathbb{Z}_{\geq 0},p\in tP\}\subset\mathbb{Z}\times M,$
graded by $\deg(t,p)=t$. Since $W$ preserves $M$ and $P$, it acts on $S$, hence on $\mathbb{Z}[S]$, yielding an action on $X_{P}$ by toric automorhpisms. The fundamental domain of the $W$-action on ×$\mathbb{R}\times V$ is ×$\mathbb{R}\times D$, so by Theorem 1.2, there is an isomorphism of graded rings
$\mathbb{Z}[t(D\cap P)]=\mathbb{Z}[(\mathbb{R}\times D)\cap S]\to\mathbb{Z}[S]^{W}.$
Applying $\operatorname{Proj}$, we obtain an isomorphism $X_{P}/W\cong X_{D\cap P}$. ∎
Example 3.5.
Suppose that $W=S_{2}$ acts on $M=\mathbb{Z}^{2}$ by coordinate permutation, and $S=\left(\mathbb{Z}_{\geq 0}\right)^{2}\subseteq M$. Then $W$ is generated by reflection across the single simple root $\alpha=(-1,1)$, with corresponding weight group $\Lambda=Z+\mathbb{Z}\cdot\lambda$, where $Z$ is the real span of $z=\frac{e_{1}+e_{2}}{2}$ and $\lambda=\frac{-e_{1}+e_{2}}{2}$. The map $\Psi:\mathbb{Z}[D\cap S]\to\mathbb{Z}[S]^{W}$ is given by
$\Psi\left(\chi^{az+b\lambda}\right)=(\chi^{z})^{a}\left(\chi^{\lambda}+\chi^{s_{\alpha}(\lambda)}\right)^{b},\quad az+b\lambda\in D\cap S.$
For instance, if $u=e_{1}+3e_{2}=4z+2\lambda$, then $\Psi(\chi^{u})=\chi^{4z}\left(\chi^{\lambda}+\chi^{s_{\alpha}(\lambda)}\right)^{2}=\chi^{4z+2\lambda}+2\chi^{4z}+\chi^{4z-2\lambda}$. See Fig. 1. Geometrically, we have that $\mathbb{A}^{2}/{S_{2}}\cong\mathbb{A}^{2}$.
$\alpha$$\lambda$$s_{\alpha}(\lambda)$$z$×$\times 1$×$\times 2$×$\times 1$
Figure 1. The lattice points in the shaded cone are $D\cap S$, and the bold lattice points marked with coefficients represent $\Psi(\chi^{u})$ for $u=e_{1}+3e_{2}$.
[…]